\underline{P365}

EXPERIMENTAL STUDY OF ATOM FLUORESCENCE LINES IN THE LASER PRODUCED PLASMA

O. A. BUKIN, \underline{E. A. SVIRIDENKOV}, Lebedev Physical Institute, 53 Leninskii pr., Moscow Russia; N. V. SUSHILOV and A. Yu. MAJOR, Pacific Oceanological Institute, Russian Academy of Sciences, 43 Baltiyskaya, Vladivostok, 690041 Russia.

Emission line shapes of laser plasma generated on the surface of an aluminium are studied experimentally. The shifts of the centers of self-reversed absorption lines with the wave length 3962 A and 3944 A are registered. Two physical mechanisms of these shifts are discussed.

Time required: 15 min Session in which paper is recomended for presentation: 11 Comment: email